OpenStudy (anonymous):

what is the integral of ((1+sqrt.x)^1/3)/(sqrt.x)

OpenStudy (anonymous):

wolframalpha will save u, use it: wolframalpha.com

OpenStudy (anonymous):

I got 2/3*(1+sqrt.x)^3/2

OpenStudy (anonymous):

click on show steps

OpenStudy (anonymous):

that not the right problem

OpenStudy (anonymous):

no i think not. if you want to cheat you have to be careful

OpenStudy (anonymous):

i dont want to cheat, i worked it out and got 2/3*(1+sqrt.x)^3/2

OpenStudy (anonymous):

rather simple u-sub yes? \[\int\frac{\sqrt[3]{1+\sqrt{x}}}{\sqrt{x}}dx\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

u=1+sqrt.x

OpenStudy (anonymous):

put \[u=1+\sqrt{x}\] \[du=\frac{1}{2\sqrt{x}}\] get \[\frac{1}{2}\int\sqrt[3]{u}du\]

OpenStudy (anonymous):

get \[\frac{1}{2}\times \frac{3}{4}u^{\frac{4}{3}}\] \[=\frac{3}{8}\sqrt[3]{(1+\sqrt{x}})^4\]

OpenStudy (anonymous):

oh no mistake!!

OpenStudy (anonymous):

sorry the constant out front should be 2, not 1/2

OpenStudy (anonymous):

\[du=\frac{1}{2\sqrt{x}}dx\] \[2du=\frac{dx}{\sqrt{x}}\]

OpenStudy (anonymous):

so instead of \[\frac{3}{8}\] it should be \[\frac{3}{2}\] out front. my mistake

OpenStudy (anonymous):

oh its raised to 4/3

OpenStudy (anonymous):

dont tell me the answer yet

OpenStudy (anonymous):

ok i let you finish it

OpenStudy (anonymous):

i got it it is :

OpenStudy (anonymous):

close but you are off by a constant.

OpenStudy (anonymous):

+ C

OpenStudy (anonymous):

should be \[\frac{3}{2}\] not \[\frac{3}{4}\]

OpenStudy (anonymous):

no i meant the constant out front. because when you do the u-sub you do not get \[du=\frac{1}{\sqrt{x}}dx\] you get \[du=\frac{1}{2\sqrt{x}}dx\]

OpenStudy (anonymous):

so you get \[2du=\frac{1}{\sqrt{x}}dx\] pull the 2 out front

OpenStudy (anonymous):

that is why it should be \[\frac{3}{2}\] not \[\frac{3}{4}\]

OpenStudy (anonymous):

but i took the derivative of the answer i got and it gave the integral

OpenStudy (anonymous):

no. the derivative of \[\sqrt{x}=\frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

so you will be off by a factor of 2

OpenStudy (anonymous):

\[\frac{d}{dx}\frac{3}{4}(1+\sqrt{x})^{\frac{4}{3}}\] \[=\frac{4}{3}\times \frac{3}{4}(1+\sqrt{x})^{\frac{1}{3}}\times \frac{1}{2\sqrt{3}}\]

OpenStudy (anonymous):

so with out the extra two out front you have the wrong answer. but easy to fix, just multiply by 2

OpenStudy (anonymous):

http://www.twiddla.com/577354

OpenStudy (anonymous):

can you cheack my work

OpenStudy (anonymous):

i am there